OpenStudy (anonymous):

jack usually mows his lawn in 5 hours. Marilyn can mow the same yard in 3 hours. how much time would it take the two of them to mow the lawn together

OpenStudy (dumbcow):

1/5 + 1/3 = 3/15 + 5/15 = 8/15 working together the work at a rate of 8/15 per hour thus it takes 15/8 hours to complete the lawn